# PADetBench: Towards Benchmarking Physical Attacks against Object Detection

### v1

@@ Appendix B Additional content of the Experiments
B.1 Genetated data for the ablation studies
@@
B.2 Detailed illustration of the gap between the attack performance of our benchmark and the reported performance in the original papers
Our benchmark addresses two important aspects that were not considered in the original papers:
1.
Strict physical dynamics: Our benchmark leverages the CARLA simulator, which provides a realistic and consistent physical environment for evaluating physical attacks on object detection models. This ensures that the attacks are evaluated under conditions that closely resemble the physical world, allowing for more accurate assessment of the robustness of the models.
2.
Evaluation in 3D space perspective: Unlike the original papers, which mainly focus on 2D evaluation, our benchmark evaluates the attack performance in a 3D space perspective. This is crucial because physical attacks can exploit vulnerabilities in the 3D structure of the objects, such as comprehensive and continuous view, scale, distance, etc. By considering the 3D space perspective, our benchmark provides a more comprehensive evaluation of the models’ robustness against physical attacks.
By addressing these two aspects, our benchmark provides a more realistic and comprehensive evaluation of the attack performance of object detection models, bridging the gap between the reported performance in the original papers and the real-world scenarios.
@@ B.4 Additional overall experimental results
Due to space constraints, we provide additional overall experimental results in this part, as shown in Table 4, 5, 6, 7, 8, 9, 10, and 11. In addition, the visualized evaluation results are shown in Fig. 9, 10, 11, 12, 13, 14, and 15.
@@ B.5 Additional ablation experiments
Due to space constraints, we provide additional overall experimental results in this part, as shown in Table 12, 13, 14, 15, 16, and 17. In addition, the visualized evaluation results are shown in Fig. 16, 17, and 18.
Appendix C Reproducibility
The reproducibility of all evaluation results in our benchmark is ensured through the provision of the necessary code in the associated Github Repository (https://github.com/JiaweiLian/Benchmarking_Physical_Attack). Furthermore, all datasets and model architectures evaluated are accessible to the public.
Appendix D License
This work is under Apache-2.0 license. More detailed information can be found in the LICENSE file in the GitHub repository: https://github.com/JiaweiLian/Benchmarking_Physical_Attack

### v3

@@ Appendix B Additional content of the Experiments
B.1 Genetated data for ablation studies
@@
B.2 A detailed illustration of the performance gap
B.2.1 Performance gap between the benchmark and the original papers
In this section, we provide an explanation for the performance gap between the reported attack performance in the original papers and the results in our benchmark. Our benchmark encompasses a wide range of physical dynamics, whereas previous validation settings are often limited to a few specific scenarios. The comprehensive physical dynamics in our benchmark reveal the shortcomings of existing object detectors and physical attacks and this is the main motivation of our work. Therefore, our benchmark might not be captured by previous validation settings, leading to the discrepancy between our results and the reviewer’s individual experiences.
In comparison, we removed various physical dynamics including weather (rain, snow, fog), lighting (nighttime), and distance (far positions), and reproduced the results of several attack methods on YOLOv7 as reported in ACTIVE [5], which are listed in Table 23. It is worth noting that these reported results are also included in our benchmark with particular evaluation settings.
Contrary to the simplified settings of these reproduced experiments, more comprehensive physical dynamics incorporated into our benchmark significantly highlight the ineffectiveness of existing physical attacks. These aspects may not have been adequately captured by previous validation settings. As illustrated in Tabel 23, when we exclude various dynamics, the effectiveness of physical attacks notably increases, thereby reducing the performance of object detectors. Therefore, our benchmark strive to encompass and align these physical dynamics for comprehensive and equitable comparisons.
B.2.2 Performance gap between attacks against vehicle and person detection
For the gap between attacks against vehicle and person detection, one reason is that these attacks are optimized to fool object detectors in particular target detection during training process. Consequently, we follow the attack purpose of the original works in this benchmark to attack specified target category accordingly for fairness, which partially accounts for the phenomenon that pedestrian detection performance is less affected regarding various attacks in comparison with car detection.
Another potential reason is that the stronger physical perturbations are optimized with consideration of 3D space and accommodate more complex physical dynamics, while physical attacks aiming to fool person detectors are commonly performed with optimized 2D patches, which work well in particular physical dynamics, as evidenced by the ablation experiments in B.2.1, which empirically demonstrate the pressing need and necessity of a comprehensive and rigorous benchmark for physical attacks.
@@ B.4 Additional overall experimental results
Due to space constraints, we provide additional overall experimental results in this part, as shown in Table 6, 7, 8, 9, 10, 11, 12, and 13. In addition, the visualized evaluation results are shown in Fig. 17, 18, 19, 20, 21, 22, and 23.
@@ B.5 Additional ablation experiments
B.5.1 Ablation study on physical dynamics
Due to space constraints, we provide additional ablation experimental results in this part, as shown in Table 14, 15, 16, 17, 18, and 19. In addition, the visualized evaluation results are shown in Fig. 24, 25, and 26.
B.5.2 Ablation study on training dataset
To further investigate the impact of the training dataset on the physical attacks, we collected ten physical attacks for fooling person detection, and the results are shown in Table 21, where the Median ASR represents the median attack success rate across the 48 detectors. It can be observed that physical attacks trained on the INRIA and COCO datasets achieve comparable performance in general.
Physical attacks | Training datasets | Median ASR |
AdvCam | ImageNet | 0 |
AdvCaT | 376 self-collected images | 0 |
MTD | - | 2 |
LAP | INRIA | 2 |
AdvPattern | Market1501 | 2 |
AdvTshirt | 40 self-collected videos | 3 |
DAP | INRIA | 5 |
NaTPatch | INRIA | 5 |
InvisCloak | COCO | 5 |
AdvTexture | INRIA | 7 |
B.5.3 Ablation study on 2D and 3D perturbations
Physical attacks that evaluate 2D adversarial patches from a frontal perspective have a significant limitation, as they do not account for the effects of multiple viewing angles in a 3D environment. Our study aims to bridge this gap by developing a comprehensive benchmark for assessing physical attacks from various angles and incorporating a broader range of physical dynamics. During our investigation, we noted a substantial drop in performance (detection rate: ) when adversarial patches were only applied to the frontal view of objects. To ensure a fair comparison and enhance the efficacy of the attacks, we expanded the application of these patches to cover the entirety of the object’s surface. Additional experiments were conducted to assess the impact of adversarial patches on frontal views using several object detection algorithms. The results are summarized in Table 22. The ’Entire Surface’ column highlights cases where the adversarial patch was applied across the entire surface of an object. The values in parentheses indicate the relative decrease in performance compared to full-surface patching.
Perturbations | Entire surface | CornerNet | VarifocalNet |
Clean | - | 87 | 80 |
Random | - | 87 | 77 |
AdvTexture | ✓ | 74 | 73 |
AdvTexture | 81(7) | 77(4) | |
AdvPatch | ✓ | 82 | 75 |
AdvPatch | 85(3) | 79(4) | |
NatPatch | ✓ | 78 | 74 |
NatPatch | 83(5) | 77(3) |
Appendix C User feedback
To ensure ease of use, we have addressed potential barriers by user feedback, such as CARLA deployment and customizing adversarial objects, by providing a comprehensive Docker installation guide for CARLA and a tutorial on customizing adversarial objects in our documentation. These resources enable users to install CARLA and customize objects in just a few minutes. We also conducted usability testing with five researchers from a well-known University and got feedback from them in the form of a survey questionnaire as shown in Table 25. The users consistently found the benchmark easy to use and provided positive feedback on its usability.
Clean | Random | CAMOU | DTA | ACTIVE | ||
YOLOv3 | Reported | 86 | 67 | 60 | 32 | 23 |
Reproduced | 86 | 66 | 62 | 33 | 23 | |
YOLOv7 | Reported | 93 | 86 | 83 | 59 | 42 |
Reproduced | 93 | 85 | 83 | 60 | 41 | |
PVT | Reported | 89 | 78 | 69 | 56 | 52 |
Reproduced | 89 | 78 | 69 | 56 | 51 |
Number | Questions |
Q1 | How easy was it to follow the Docker installation guide for CARLA? (Rating 1-5) |
Q2 | How helpful was the tutorial on customizing adversarial objects in the documentation? (Rating 1-5) |
Q3 | Were you able to successfully deploy CARLA using the provided resources? (Yes or No) |
Q4 | Were you able to successfully customize adversarial objects using the provided resources? (Yes or No) |
Q5 | Overall, how satisfied are you with the ease of CARLA deployment and customizing adversarial objects? (Rating 1-5) |
Questions | User1 | User2 | User3 | User4 | User5 |
Q1 | 4 | 5 | 5 | 4 | 5 |
Q2 | 5 | 5 | 5 | 5 | 5 |
Q3 | Yes | Yes | Yes | Yes | Yes |
Q4 | Yes | Yes | Yes | Yes | Yes |
Q5 | 4 | 5 | 5 | 4.5 | 5 |